2. The sheiving Was sond During the current year, Ramirez Developers disposed of plant assets in the following transactios LEM 9.4A sal of Plant Office equipment costing $26,000 was given to a scrap dealer at no charge. At the date of disposal, accumulated depreciation on the office equipment amounted to $25,800 Feb. 10 Ramirez sold land and a building to Claypool Associates for $900,000, receiving $100,000 cash and a five-year, 9 percent note receivable for the remaining balance. Ramirezs records showed the following amounts: Land, $50,000; Building, $550,00 Accumulated Depreciation: Building (at the date of disposal), $250,000. Apr. 1 Ramirez traded in an old truck for a new one. The old truck had cost $26,000, and its accumulated depreciation amounted to $18,000. The list price of the new truck was $39,000, but Ramirez received a $10,000 trade-in allowance for the old truck and paid only $29,000 in cash. Ramirez includes trucks in its Vehicles account. Aug. 15 Oct. 1 Ramirez traded in its old computer system as part of the purchase of a new system old system had cost $15,000, and its accumulated depreciation amounted to SIT The new computers list price was $8,000. Ramirez accepted a trade-in allowanc of $500 for the old computer system, paying $1,500 down in cash and issuing a o 8 percent note payable for the $6.000 balance owed. Instructions L04 PRO a. Prepare journal entries to record each of the disposal transactions. Assume that deprecid expense on each asset has been recorded up to the date of disposal. Thus, you need n Altern Depre the accumulated depreciation figures stated in the problem. income statement? Explain. cial reporting of unrealized gains and losses on marketable securities discu b. Will he gains and losses recorded in part a above affect the gros profi reported gros profit reported in Ramine e. Explain how the financial reporting of gains and losses on plant assets differs fron chape rom the finat PROBLEM 9.5A
Answer #1

Answer to Part (a)

Feb 10

Here the net book value of office equipment is = 26,000 - 25,800 = 200, this office equipment is being discarded for nothing so this 200 is the loss the journal entries will be as follows

(a)

Asset Disposal account debit 26,000

Office equipment account credit 26,000

(b)

Accumulated Depreciation on Office Equipment debit 25,800

Loss on disposal of office equipment debit 200

Asset Disposal credit 26,000

April 1

Here net book value of land and building is = 50,000 + 550,000 - 250,000 = 350,000

The land and building is being sold for 900,000 the difference (900,000 - 350,000 = 550,000) this is the gain on sale of land and building

(a)

Asset disposal account debit 600,000

Land account credit 50,000

Building account credit 550,000

(b)

Accumulated depreciation debit 250,000

Asset disposal account credit 250,000

(c)

Cash account debit 100,000

Notes receivables debit 800,000

Asset disposal account credit 350,000

Gain on disposal of asset 550,000

August 15

Here net book value of old truck is = 26,000 -18,000 = 8,000 and for this a trade in discount of 10,000 has been received so the difference of 2,000 is the gain.

(a)

Vehicle account debit 39,000

Cash account credit 29,000

Asset disposal account credit 10,000

(b)

Asset disposal account debit 26,000

Vehicle account credit 26,000

(c)

Accumulated depreciation on vehicle debit 18,000

Asset disposal account credit 18,000

(d)

Asset disposal account debit 2,000

Gain on sale of asset 2,000

October 1

Here net book value of old computer = 15,000 - 11,000 = 4,000 for this a trade in discount of 500 has been received. the difference i.e. 3,500 is loss on disposal of computer.

(a)

Asset disposal account debit 15,000

Computer account credit 15,000

(b) Accumulated depreciation on computer debit 11,000

Asset disposal account credit 11,000

(c)

Computer account debit 8,000

cash account credit 1,500

Notes payable credit 6,000

Asset disposal account 500

(d)

Loss on sale of asset debit 3,500

Asset disposal account credit 3,500

Answer to Part b

The gain or loss recorded in part (a) will not affect the gross profit as this gain or loss is transferred to profit and loss account and it will affect the net profit and not the gross profit.

Answer to Part C

Situation one if no gain or loss had been recognized earlier on account of fair value accounting

The gain or loss on sale of assets is transferred to profit and loss account.

But in case of unrealized gain or loss on marketable securities the treatment is different it is as follows

(a) first the realized gain or loss is adjusted against the existing balance of unrealized amount of gain or loss recognized earlier.

(b) Any balance thereafter is transferred to profit and loss account.